### v1

@@ 4.2 Metrics
4.3 Implementation Details
X-SG2S is implemented with PyTorch, along with an off-the-shelf 3DGS render implemented in CUDA. All models are trained on single A6000ada with the Adam optimizer.
@@ 4.3 Implementation Details
We choose binary message of 8bit×$\times$128 as 1D watermark and 3DGS object with 10000 points as 3D watermark. For 2D watermark, we embed them via a pretrained AE encoder then split them in order to get 16$\times$32768 message patches. In the testing phase, we reconstruct the information fragments into a predetermined shape and then pass them through the AE decoder to obtain the image. And the number of SH parameters is 75 for each GS point.
4.4 Main Results
4.4.1 ablation experiments
Figure 4: Interaction Blocks: a) “Multiply”: it multiplies the self-score and the cross-score to generate final-score. b) “Add”: it uses efficient cross attention to generate cross-feature and set self attention to generate self-feature. Then it concatenates them and calculates final-score. c) “Concatenate”: it concatenates cross-feature and self-feature together and calculates final-score. d) “Normal”: it just uses 3DGS original points as input and passes them through n layers of MLP gate. e) “Random”: it randomly selects k locations to add extra message. It can not be learned.
@@ 4.4 Main Results
Method Multiply Add Concatenate Normal Random
PSNR(org)$\uparrow$ 25.341 23.81 24.271 23.857 24.406
SSIM(org)$\uparrow$ 0.848 0.829 0.833 0.833 0.759
LPIPS(org)$\downarrow$ 0.178 0.214 0.208 0.203 0.246
$\uparrow$ 1.0 1.0 0.814 1.0 0.564
PSNR(image)$\uparrow$ 36.0062 36.0426 36.0877 35.8784 3.9009
PSNR(3D obj)$\uparrow$ 22.1400 23.7939 22.5261 22.1902 22.9520
SSIM(3D obj)$\uparrow$ 0.8603 0.8762 0.8653 0.8606 0.8728
LPIPS(3D obj)$\downarrow$ 0.1444 0.1387 0.1491 0.1555 0.1626
Table 1: Multimodal joint training with different structures of interaction block.
@@
PSNR(org)$\uparrow$ 27.821 27.4 26.398 28.25
SSIM(org)$\uparrow$ 0.871 0.866 0.858 0.843
LPIPS(org)$\downarrow$ 0.125 0.139 0.153 0.144
@@
PSNR(image)$\uparrow$ / 35.879 / /
PSNR(3D obj)$\uparrow$ / / 22.28 28.807
SSIM(3D obj)$\uparrow$ / / 0.869 0.942
LPIPS(3D obj)$\downarrow$ / / 0.154 0.067
Table 2: Single-modality watermark injection and extraction training.
Method X-SG2S(Joint) X-SG2S(Individual)
PSNR(org)$\uparrow$ 27.788(1D) / 26.125(2D) / 26.493(3D) 27.821(1D) / 27.4(2D) / 26.398(3D)
SSIM(org)$\uparrow$ 0.871(1D) / 0.859(2D) / 0.860(3D) 0.871(1D) / 0.866(2D) / 0.858(3D)
LPIPS(org)$\downarrow$ 0.125(1D) / 0.154(2D) / 0.149(3D) 0.125(1D) / 0.139(2D) / 0.153(3D)
@@
PSNR(image)$\uparrow$ 36.0085 35.879
PSNR(3D obj)$\uparrow$ 22.2160 22.28
SSIM(3D obj)$\uparrow$ 0.8608 0.869
LPIPS(3D obj)$\downarrow$ 0.1443 0.154
Table 3: Comparison between adding a single watermark under joint training and individual training.
4.4.2 single modality injection test
We also test X-SG2S’s performance on training single modality. In Table 2, we can see that X-SG2S can well adapt the task of 1 to 3D watermarking. To our delight, we discovered that the scene with watermark outperformed the original in terms of both SSIM and LPIPS metrics. Meanwhile, in Table 3, we compare the differences between adding a single watermark under joint training and under individual training. We can see that two results are almost identical. This demonstrates the effectiveness of our method. The image results are all presented in the appendix.
4.4.3 The Exploration of Loss Function
Method without optional losses with optional losses
PSNR(org)$\uparrow$ 24.148 25.341
SSIM(org)$\uparrow$ 0.840 0.848
LPIPS(org)$\downarrow$ 0.182 0.178
$\uparrow$ 1.0 1.0
PSNR(image)$\uparrow$ 35.8708 36.0062
PSNR(3D obj)$\uparrow$ 21.9596 22.1400
SSIM(3D obj)$\uparrow$ 0.8568 0.8603
LPIPS(3D obj)$\downarrow$ 0.1564 0.1444
Table 4: Comparison with and without optional loss functions.
@@ 4.4.3 The Exploration of Loss Function
4.4.4 The Exploration of Text Watermarking
We surprisingly found that X-SG2S can achieve nearly 100% accuracy on text watermark. So, we have delved deeper into its exploration. We conduct two experiments: 1) Investigate the impact of binary bit size on watermarking. 2) Investigate the impact of the sentence’s length on watermarking.
For 1), we test the bit sizes of 8, 16, 24, 32, 40 which are the most commonly used, with its length of 12288. From Figure 5, we can see that under our experimental conditions, the size of 16 can achieve the best result. And the experiment also shows that the bit sizes does not have a great impact on the original scene.
For 2), we test the lengths of 128, 4096, 8192, 12288, 16384, 20480 with its size of 8. From Figure 5, we can see that as the length increases, the integrity of the original scene tends to decline, which implies that the length of the message patches affects the completeness of the original scene.
Figure 5: The above row of images shows the experiment 1) and the below row shows 2). It shows the original measurements(triangle) and the trend(line) when X-SG2S achieves nearly 100% accuracy. The red one is the metric of the original scene without watermark.
5 Conclusion
@@
Appendix A Appendix
Figure 6: This image demonstrates its capability to inject and extract watermarks across all three modalities simultaneously. And the X-SG2S was jointly trained with three modalities’ watermarks. The first column is the restored logo and the second is ground truth. The third one is the restore objects, the forth is the image render by dataset, the fifth column is the ground truth. The sixth column is the scene we added message and the last column is the ground truth. We can see that the results are quite impressive. And binary precision is nearly 1.0.

### v2

@@ 4.2. Metrics
4.3. Implementation Details
X-SG2S is implemented with PyTorch, along with an off-the-shelf 3DGS render implemented in CUDA. All models are trained on single RTX4090 with the Adam optimizer.
@@ 4.3. Implementation Details
We choose binary message of 48bit×$\times$1024 as 1D watermark and 3DGS object with 10000 points as 3D watermark. For 2D watermark, we use feature sparse DCAE to get 16$\times$2048 message patches. And the number of higher order SH parameters is 48 for each GS point.
4.4. Main Results
4.4.1. ablation experiments
Figure 6. Interaction Blocks: a) “Multiplication”: it multiplies the self-score and the cross-score to generate final-score. b) “Addition”: it uses efficient cross attention to generate cross-feature and set self attention to generate self-feature. Then it concatenates them and calculates final-score. c) “Concatenation”: it concatenates cross-feature and self-feature together and calculates final-score. d) “Normal”: it just uses 3DGS original points as input and passes them through n layers of MLP gate. e) “Random”: it randomly selects k locations to add extra message. It can not be learned.
@@ 4.4. Main Results
Method Multiplication Addition Concatenation Normal Random
PSNR(org)$\uparrow$ 27.666 25.812 25.261 25.951 25.614
SSIM(org)$\uparrow$ 0.868 0.839 0.833 0.845 0.759
LPIPS(org)$\downarrow$ 0.135 0.214 0.208 0.203 0.246
$\uparrow$ 1.0 1.0 1.0 1.0 1.0
PSNR(image)$\uparrow$ 30.537 30.421 29.991 30.174 30.552
PSNR(3D obj)$\uparrow$ 21.140 21.794 21.201 21.141 21.231
SSIM(3D obj)$\uparrow$ 0.843 0.846 0.855 0.842 0.848
LPIPS(3D obj)$\downarrow$ 0.166 0.159 0.163 0.157 0.158
Table 1. Multimodal joint training with different structures of interaction block.
@@
PSNR(org)$\uparrow$ 27.837 27.811 27.759 28.25
SSIM(org)$\uparrow$ 0.871 0.866 0.870 0.843
LPIPS(org)$\downarrow$ 0.125 0.131 0.133 0.144
@@
PSNR(image)$\uparrow$ / 30.537 / /
PSNR(3D obj)$\uparrow$ / / 21.144 28.807
SSIM(3D obj)$\uparrow$ / / 0.843 0.942
LPIPS(3D obj)$\downarrow$ / / 0.165 0.067
Table 2. Single-modality watermark injection and extraction.
4.4.2. single modality injection test
In Table 2, we test our X-SG2S on adding sigle watermark. We can see that X-SG2S can well adapt the task of 1 to 3D watermarking. We can see that the performance of watermark extraction are almost identical. This demonstrates the effectiveness of our method.
4.4.3. The Exploration of Loss Function
Method without optional losses with optional losses
PSNR(org)$\uparrow$ 25.241 27.666
SSIM(org)$\uparrow$ 0.842 0.868
LPIPS(org)$\downarrow$ 0.172 0.135
@@ 4.4.3. The Exploration of Loss Function
PSNR(image)$\uparrow$ 29.132 30.537
PSNR(3D obj)$\uparrow$ 20.616 21.140
SSIM(3D obj)$\uparrow$ 0.837 0.843
LPIPS(3D obj)$\downarrow$ 0.166 0.166
Table 3. Comparison with and without optional loss functions.
@@
4.5. Fidelity
According to Figure 7, we show the importance of SH. The left column shows X-SG2S injects the watermark to all the SH while the right shows the watermarks are only injected to the higher order SH. It seems that only inject watermarks to higher order SH can achieve good results. From the visual perspective, the invisible watermark cannot be detected by the naked eye which means we protect the original scene well and hide the watermarks well. Meanwhile,we have drawn the residual map of the invisible watermark in Figure 8. It can be observed that our method can effectively hide the watermark without detection. Figure 9 shows that our X-SG2S can inject and extract watermarks well.
Figure 7. This image illustrates the different effects of watermarking all SH coefficients versus watermarking only the high-order SH coefficients.
Figure 8. This image demonstrates the effectiveness of our approach. The first column shows the original scene images rendered by MVSplat. The second column displays the images after watermark embedding. The third column presents the 10×residuals between the two images. It can be seen that our model is capable of effectively concealing the watermark.
Figure 9. This image demonstrates the watermarks. The first column shows the original watermarks provide by Logo-2K. The second column displays the watermarks extracted by X-SG2S. The third column presents the GS objects extracted by X-SG2S. The final column displays the GS objects provided by Gamba.
4.6. Robustness
To evaluate the robustness of X-SG2S, we have subjected the Gaussians to degradation using random pruning methods. Quantitative metrics are shown in Tab 4. The results indicate that our method effectively withstands the degradation process.
Pruning Rate PSNR(org) SSIM(org) LPIPS(org) precision PSNR(image) PSNR(3D obj) SSIM(3D obj) LPIPS(3D obj)
5% 26.799 0.821 0.246 1.0 30.511 20.568 0.843 0.169
10% 25.784 0.775 0.294 1.0 29.847 20.572 0.842 0.170
15% 24.661 0.728 0.333 1.0 29.931 20.520 0.842 0.171
20% 23.485 0.682 0.366 1.0 29.620 20.484 0.842 0.173
25% 22.273 0.635 0.398 1.0 29.571 20.433 0.841 0.174
Table 4. Robustness testing by random pruning GS points.
PSNR(org) SSIM(org) LPIPS(org) precision PSNR(image) PSNR(3D obj) SSIM(3D obj) LPIPS(3D obj)
GS without wm 28.25 0.843 0.144 0.395 7.719 11.934 0.835 0.319
Table 5. To verify whether the detection model can distinguish if the original model contains a watermark.
4.7. Precise Identification
In the real case, we cannot indiscriminately detect watermarks. That is to say, we cannot falsely identify a watermark in a model that does not have one. To illustrate this, we randomly select some GS scenes without watermarks and observe that our model fails to detect any watermark. Table 5 demonstrates that our model can effectively reduce false positives without the need for an additional discriminator.
5. Conclusion